# Entropy-conservative high-order methods for high-enthalpy gas flows

### v1

@@ 4.1.2 Infinite harmonic oscillator
Next, we compare the energy flux (39) to the flux derived in [34] for the case of the infinite harmonic oscillator model. The gas was assumed to be molecular oxygen (with a molecular mass $m=5.3134\times 10^{-26}$ kg), and a characteristic vibrational temperature of $\theta_{v}=2273.5$ K was used. We consider the following three scenarios:
  1. 1.
    Comparison of the energy flux (39) to the flux derived in [34] assuming the temperatures are known exactly (no impact of error in the temperature estimation).
  2. 2.
    Comparison of the energy flux (39) to the flux derived in [34] using linear interpolation to infer the temperature from the internal energy for use in (39), but assuming the temperature to be known exactly for the flux from [34], thus treating it as an exact benchmark.
  3. 3.
    Comparison of the flux derived in [34] with an exactly known temperature to its version where the temperature is computed from the internal energy using the Mutation++ library [41].
@@
Figure 3: Relative error in the energy flux as a function of the discretization step $\Delta T$, infinite harmonic oscillator. Different line styles correspond to different temperature jumps, different line colors correspond to the different scenarios described above.
Figure 3 shows the relative errors in the energy flux (compared to the flux from [34] assuming exactly known temperatures) for different temperature jumps as a function of the discretization step $\Delta T$. The black lines correspond to scenario 1: exact temperatures are given in the flux functions, and the error is only due to approximation of jumps in the entropy and energy. The red lines correspond to scenario 2: linear interpolation is used to compute the temperature from the energy for use in Eqn. (39). Finally, the blue lines correspond to the error in the flux from [34] arising due to the inference of temperature from the energy using the Mutation++ library. We see that at values of $\Delta T$ up to 500 K, the error in the flux is strongly affected by the error due to the inexact estimation of temperature from the energy, at higher values of $\Delta T$, the error in the estimation of the jumps in energy and entropy becomes dominant. Of course, in practice, small discretization values are usually used (1-5 K), in which case the relative error in the flux when comparing it to the exact flux with exactly known temperatures is approximately 10-6–10-5. When one compares the flux from [34], where temperatures where computed using Mutation++, to its exact version (with no error in the temperature), the relative error is 10-5. As Mutation++ uses a non-linear solver to deduce the temperature, the error is independent of $\Delta T$. Thus, at small values of $\Delta T$, the proposed flux expression (39) achieves excellent error despite using linear interpolation both for estimation of temperature and estimation of the jumps in internal energy and entropy.
@@ 4.1.3 Computational cost comparison
We carry out a brief comparison of the computational cost of calculating the flux vector (36)–(39) and the computational cost of calculating the flux vector as given in [34]. Both fluxes were implemented in the Julia programming language. On a i9-13900K processor the computation of the flux vector using Eqns. (36)–(39), including the estimation of temperature from the internal energy using linear interpolation, required 17.8 ns on average, whereas computation of the flux vector as given in [34] without accounting for the cost of inverting the energy to obtain temperature required 26.8 ns on average, approximately 50% more computational time. Estimation of temperatures of the left and right states from the internal energy using the C++ interface of the Mutation++ (compiled using the -O3 compiler flag) library adds another 56 ns of CPU time to each flux evaluation, thus bringing the total time required to compute the flux vector derived in Peyvan et al. [34] to 82.8 ns on average.
For a fairer comparison, the linear interpolation of temperature from energy was replaced by a non-linear Newton-Raphson solver using the same relative tolerance of $10^{-12}$ as Mutation++, as well as the same starting guess for the temperature $T$. In that case, evaluation of the flux vector proposed in the present work required 13 ns on average, as the temperature estimation is now performed separately. The computation of the temperatures of the left and right states from the internal energy required 53 ns on average. Thus, the total cost of evaluating the flux proposed in the present work when a non-linear solver is used to infer the temperatures is on average 66 ns. The flux of Peyvan et al. [34] in conjunction with the Mutation++ library requires approximately 83 ns per evaluation, a 25% increase in computational cost. When the temperature inversion costs are not accounted for, the evaluation of the flux of Peyvan et al. [34] is approximately twice as slow than the computation of the flux (36)–(39).
As a baseline comparison, the flux of Chandrashekar [7] for a calorically perfect gas requires 12 ns per flux evaluation, as the temperature can be computed from the energy via simple analytical inversion, and no look-ups and interpolations are required. Computation of the volume flux is one of the most expensive parts of the solver, with the current implementation in Trixi.jl spending between 50% and 75% of the total run-time computing the volume flux (depending on the polynomial order used in the simulation, with higher orders requiring more flux evaluations). As such, reducing the computational cost of the flux function is crucial to the efficiency of the solver [37]. The flux (36)–(39) derived in the present work thus has been shown have excellent error and computational performance. Moreover, it is important to emphasize again that its performance is independent of the model chosen for the description of the internal energy of the gas, as exact evaluations are performed only during the pre-processing step. It can also be seen that the inversion of energy in order to obtain the temperature comprises the by far most expensive part of the flux evaluation (up to 80% of the total time), although this can be somewhat improved by supplying a good initial guess for the temperature, reducing the number of iterations until convergence.
@@ 4.2 Periodic flow
To assess the performance of the developed flux, we consider a two-dimensional test case with periodic boundary conditions, similar to the one investigated in [34]. A square domain $[0,1]\times[0,1]$ was assumed. The gas was taken to be molecular nitrogen, with a molecular mass of $m=4.6517\times 10^{-26}$ kg and with the vibrational spectrum modelled by an infinite harmonic oscillator model with $\theta_{v}$ = 3393.5 K. The temperature discretization $\Delta T$ was taken to be 1 K. The pressure in the domain was taken to be constant as $p=195256$ Pa, and a velocity vector $\mathbf{v}=(11450;0)^{\mathrm{T}}$ m/s was assumed. A temperature profile varying in the $x$ direction was prescribed as $T(x)=9000.0+200\sin(2\pi x)$. A uniform $64\times 64$ grid was used for the discretization of the domain, and 3-rd degree polynomials were used in the DG method. All shock-capturing functions and positivity-preserving limiters were turned off, and the flux (36)–(39) was also used for the interface fluxes (instead of a Riemann solver), to ensure that no stabilization or dissipation schemes can impact the entropy conservation. The total entropy production rate (over all cells in the simulation domain), which can be defined as
$s_{t,\mathrm{tot}}=\sum\frac{\partial s}{\partial\mathbf{u}}\cdot\frac{\partial\mathbf{u}}{\partial t}$ (46)
was computed via analysis callbacks available in Trixi.jl.
@@
Figure 4 shows the total entropy production rate plotted over time. It can be clearly seen that the entropy is not conserved to machine precision, due to the aforementioned errors in the computation of temperature and jumps in the energy and integral part of the entropy; however, the overall magnitude of the entropy production rate is still very small.
@@ 4.3 Blast wave
Here $\phi=\tan^{-1}(y/x)$. The initial conditions of the problem correspond to a circular-shaped region with a radius of 0.5 m of higher density and pressure with a constant outward radial velocity. As in the previous case, the gas was assumed to be molecular nitrogen, with a vibrational spectrum described by the infinite harmonic oscillator model. A square domain of size $[-2,2]\times[-2,2]$ was used, discretized by a uniform $64\times 64$ grid. 3-rd degree polynomials were used in the DG method.
@@
Figure 7 shows the total entropy production rate plotted over time. Similar to the previous test case, it can be seen that whilst there is an error in the entropy production rate that is larger than machine precision, it does not increase over time and is still of a very small magnitude.
@@ 4.4 Flow around a cylinder
Finally, we consider a two-dimensional flow around a cylinder. The free-stream parameters were chosen to be $v_{\infty}=5956$ m/s, $p_{\infty}=476$ Pa, $T_{\infty}=901$ K, based on the experimental HEG cylinder test-case [22]; the cylinder diameter is 0.045 m. The gas was assumed to be molecular oxygen with a molecular mass $m=5.3134\times 10^{-26}$ kg. The flow speed is approximately Mach 10.5, depending on the chosen model for the internal energy. Two models for the internal energy were considered:
@@
    Infinite harmonic oscillator with $\theta_{v}=2273.5$ K.
@@
    Cut-off anharmonic oscillator with $\theta_{v}=2273.5$ K and $\theta_{v,\mathrm{anh}}=17.366$ K. The cut-off was performed not based on the dissociation energy of oxygen, but based on the level number where the vibrational energy begins to decrease with increasing vibrational level (as it is given by a quadratic dependence on the level number). This corresponds to 66 vibrational levels accounted for.
@@
The simulations were performed on a simple shock-fitted structured $60\times 60$ grid. A single cubic curve was fitted to the shock shape. As it does not fit the shock perfectly, some minor errors in the density and pressure are expected near certain regions of the shock, where the mesh and the shock are not well-aligned. Figure 8 shows an example of the simulation grid, specifically, the grid used for simulations with the infinite harmonic oscillator model. The origin of the cylinder is located at $(x,y)=(0,0)$. In the analysis of the simulations presented further below, flow quantities are considered along the stagnation line, given by $y=0$, and the outflow boundary, given by $x=0$. The following boundary conditions were applied in the simulation: supersonic inflow at the inflow boundary (highlighted in purple), supersonic outflow for the outflow boundary (highlighted in orange), and slip conditions at the stagnation line (highlighted in red) and cylinder surface (highlighted in blue).
The developed entropy-conserving flux was used for the volume fluxes, and the local Lax-Friedrichs flux was used for the surface fluxes in order to provide sufficient numerical diffusion for the strong shocks present in the flow. The sub-cell shock capturing method of Hennemann et al. [16] was used, along with the positivity-preserving limiter of Zhang and Shu [46].
The DLR TAU solver [28, 14] was used as a benchmark solver for this case with a similar grid and parameters governing the gas properties. The AUSM+ flux [27] was used, along with local time-stepping and a first-order implicit Backward Euler solver. In the case of the infinite harmonic oscillator and cut-off anharmonic oscillators the input thermodynamic data files for DLR TAU were adjusted to provide agreement between the thermodynamic data implemented in Trixi and the data used in the TAU code. Yet there remained some discrepancies in the specific heats. Therefore, to first verify the basic solver components (discretization, shock-capturing, limiting, etc.) for the considered flow, the simplified case with a constant specific heat was considered. The gas was assumed to have the molecular mass of nitrogen, and a heat capacity ratio $\gamma=1.4$ was assumed.
Figure 9: Flow-field pressure computed using the TAU code (left) and the developed entropy-conservative approach implemented in Trixi.jl (right) for the case of a constant specific heat ($\gamma=1.4$).
Figure 9 shows the computed flow-field pressure for TAU and Trixi.jl simulations grid, with the Trixi.jl results computed using 2-nd order polynomials. Excellent agreement can be observed between the solvers. Since the Trixi.jl solution used a grid with several larger cells in the pre-shock region, the visualized results differ somewhat due to the different domains. For a more quantitative comparison, we evaluate the flow properties along the stagnation line.
Refer to captionRefer to caption
@@
Figure 11: Flow-field pressure computed using the TAU code (left) and the developed entropy-conservative approach implemented in Trixi.jl (right) for the case of the infinite harmonic oscillator.
Now we consider the cases with temperature-dependent specific heats. Figure 11 shows the computed flow-field pressure for the case of the vibrational spectrum modelled by an infinite harmonic oscillator. Excellent qualitative agreement can be seen between the simulations: even the smeared pressure jump region near the shock (where the shock is strong and not well-aligned with grid) is reproduced by the Trixi.jl implementation.
Refer to captionRefer to caption
@@
For a more detailed analysis, we look at the flow-field properties along the stagnation line $y=0$ (assuming that the center of the cylinder is located at $(0,0)$). Figure 12 shows the pressure and temperature along the stagnation line (the surface of the cylinder is located at $x=-4.5$ cm) computed using TAU and Trixi.jl for the two different vibrational spectrum models, i.e. infinite harmonic oscillator and the cut-off anharmonic oscillator, with the cut-off determined by the vibrational level where the vibrational energy starts decreasing with increasing level number. The Trixi.jl simulation used 3-rd order polynomials. We see good agreement between the solutions, although the Trixi.jl simulation exhibits a slightly higher post-shock temperature and a larger shock stand-off distance in the case of the infinite harmonic oscillator, and a lower post-shock temperature and smaller stand-off distance in the case of the cut-off anharmonic oscillator. To investigate the source of the discrepancy, during the TAU simulations, the computed ratio of specific heats $\gamma$ was output along with the other flow quantities. We then can plot it as a function of temperature in the corresponding cell, allowing to investigate the temperature dependence of $\gamma$.
Figure 13: The heat capacity ratio $\gamma$ for molecular oxygen as a function of temperature computed using TAU (red lines) and the various models considered in the present work (black lines).
Figure 13 shows the heat capacity ratio of molecular oxygen computed using TAU as a function of temperature for two models for the vibrational spectrum: the infinite harmonic oscillator and the cut-off anharmonic oscillator. It also shows the values of $\gamma$ computed using the different vibrational spectra models used in the present work, already shown on Fig. 1. It can be seen that for the infinite harmonic oscillator model the TAU simulations had lower values of $\gamma$ than the implementation in the present work, especially at higher temperatures, which explains the smaller stand-off distance computed using the TAU code [42]. Conversely, for the cut-off anharmonic oscillator model the results produced by TAU had higher values of $\gamma$ in most of the considered temperature range, leading to a larger stand-off distance. Therefore, it can be concluded that the differences observed in the flow-field are due to differences in the implementation of thermodynamic gas properties.
Finally, we analyse the flow quantities in the expansion region, namely, along the outflow boundary given by $x=0$. As the temperatures are lower in that region, and the shock is significantly weaker, the impact of discrepancies in the thermodynamical properties is expected to be less significant. Slightly more differences were observed between Trixi.jl simulations using different grid and polynomial orders in this region, therefore, more results are presented: simulations were carried out for 2-nd, 3-rd, and 4-th order polynomials on $30\times 30$ and $60\times 60$ grids.
Refer to captionRefer to caption
@@
Figure 14 shows the pressure and temperature in the expansion region along the line $x=0$ (the surface of the cylinder is located at $y=4.5$ cm) for the infinite harmonic oscillator model of the vibrational spectrum. Solid lines correspond to the $60\times 60$ grid, dotted lines correspond to the $30\times 30$ grid. We see that all simulation results agree very well, with only the 2-nd order DG simulation on the coarse $30\times 30$ grid deviating slightly in terms of pressure. The Trixi.jl-based DG simulations predict a sharper shock, but use of the higher-order polynomials also leads to the presence of irregularities (seen on the temperature profile at $y=8$ cm).
Refer to captionRefer to caption

### v2

@@ 4.1.2 Infinite harmonic oscillator
Next, we compare the energy flux (41) to the flux derived in [42] for the case of the infinite harmonic oscillator model. The gas was assumed to be molecular oxygen (with a molecular mass $m=5.3134\times 10^{-26}$ kg), and a characteristic vibrational temperature of $\theta_{v}=2273.5$ K was used. The same flow conditions and temperature jumps were considered as for the previous case, and values of the tolerance $r$ was again chosen to be $\Delta T/2$. For the evaluation of both fluxes, the temperatures were assumed to be known exactly, i.e. no impact of the non-linear solver was considered. It should be noted that the numerical flux proposed in [42] relies upon a Taylor expansion of a hyperbolic sine function of a non-linear combination of $T_{-}$ and $T_{+}$, using only the first four terms of the series. For larger temperature jumps, the expansion was found to require one extra term to provide a sufficiently accurate value.
@@
Figure 3: Relative error in the energy flux as a function of the discretization step $\Delta T$, infinite harmonic oscillator.
Figure 3 shows the relative errors in the energy flux (compared to the flux from [42] assuming exactly known temperatures) for different temperature jumps as a function of the discretization step $\Delta T$. The black lines show the error relative to the flux of Peyvan et al. [42] when a 5-term Taylor expansion is used, the red line shows the error relative to the same flux when only 4 terms are used for the larges temperature ratio $T_{+}/T_{-}=10$.
Compared to the case of the calorically perfect gas, the choice of temperature discretization step $\Delta T$ has a noticeable impact on the accuracy of the flux, as it affects the accuracy of the computation of $c_{v}(T)$ and as a result, the accuracy of the computation of $\eta(T)$.
Again, it can be seen that at small values of $\Delta T$, the proposed flux expression (41) achieves excellent error. The error in this case is specific to the linear interpolation used in the present work, however, the developed numerical flux formulation allows for use of higher-order interpolations or even exact expressions for energy, specific heats, and the integral part of the entropy. A practical large-scale simulation will also require maintaining a balance between computational cost and accuracy — as will be discussed in the next subsection, use of simplified interpolations based on tabulated values can provide a noticeable speed-up compared to use of exact analytical expressions. Even at extremely large values of $\Delta T$, i.e. 100–1000 K, the error compared to the exact flux derived in [42] is still less than 1%.
It can also be seen that for the largest temperature jump considered, the 4-term expansion used in [42] proves to be insufficiently accurate. In fact, the number of terms required in the Taylor expansion used can be considered to be an implicit numerical parameter, higher values of which may lead to an increase in accuracy, but also to an increased computational cost. Whereas in the present work the main parameter is the explicitly defined $\Delta T$, which does not affect the computational cost apart from the pre-processing step at the start of the simulation.
@@ 4.1.3 Computational cost comparison
We carry out a brief comparison between the computational cost of calculating the flux vector (38)–(41) in conjunction with Eqn. 46, the computational cost of calculating the flux vector as given in [42], and the computational cost of evaluating the flux of Chandrashekar [9] for a calorically perfect gas. The flux of Chandrashekar [9] was included in the analysis as a baseline, despite its assumption of a calorically perfect gas. All fluxes were implemented in the Julia programming language.
Present work Peyvan et al. [42], 4-term expansion Chandrashekar [9]
17 ns 23 ns 11 ns
Table 1: Time required for one flux evaluation, omitting computation of temperature from energy.
Table 1 presents the computational costs of one flux evaluation for the three flux functions considered. As expected the flux of Chandrashekar [9] is computationally the most of efficient, since for the calorically perfect gas no interpolations are required. For the flux derived in  [42], the original proposed 4-term Taylor expansion of the hyperbolic sine function was used; as discussed above, for large temperature jumps, this may be insufficient and more terms may need to be included, leading to higher computational costs. It can be seen that the proposed flux function is approximately 50% slower than the flux of Chandrashekar [9], but still roughly 25% faster than the analytical flux for the infinite harmonic oscillator model with the 4-term Taylor expansion [42].
Inversion using interpolated $\varepsilon_{\mathrm{int}}$, $c_{v}$ Inversion using exact formulas for energy and $c_{v}$
75-85 ns 400 ns
Table 2: Time required to compute $T$ from internal energy.
As already mentioned, the computational costs of the flux (38)–(41) and the flux from [42] do not include the cost of the inversion of the energy in order to obtain the temperature. Assuming that $T_{-}=1000$ K and $T_{+}=1500$ K, and using an initial guess of $300$ K as the starting solution, Mutation++ (compiled using the -O3 compiler flag) requires approximately 400 ns to compute $T_{-}$ and $T_{+}$ from the corresponding energies; it does so using exact analytical formulas for the vibrational energy and specific heat of vibrational degrees of freedom, as given by (5). Using the piece-wise linear interpolation of the internal energy and specific heat, an implementation of Newton’s method in Julia requires approximately 75-85 ns for the computation of $T_{-}$ and $T_{+}$ when also using an initial guess of $300$ K and the same relative and absolute tolerance of $10^{-12}$. For a more complicated internal energy function (such as (2) the cost of inverting the energy to obtain temperature will be even higher. These results are summarized in Table 2.
Therefore, it can be concluded that the flux function itself (38)–(41) is computationally very efficient, especially for cases of a calorically imperfect gas, as it allows for use of arbitrary internal energy functions without the computational complexity of those having an impact on the cost of the flux evaluations. The use of a piece-wise linear approximation for the internal energy and specific heats also leads to a significant speed-up of the Newton solver used to compute the temperature from the internal energy.
@@ 4.2 Periodic flow
To assess the performance of the developed flux, we consider a two-dimensional test case with periodic boundary conditions, similar to the one investigated in [42], but with a larger temperature variation. A square domain $[0,1]\times[0,1]$ was assumed. The gas was taken to be molecular nitrogen, with a molecular mass of $m=4.6517\times 10^{-26}$ kg and with the vibrational spectrum modelled by an infinite harmonic oscillator model with $\theta_{v}$ = 3393.5 K. The temperature discretization $\Delta T$ was taken to be 1 K. The pressure in the domain was taken to be constant as $p=195256$ Pa, and a velocity vector $\mathbf{v}=(11450;0)^{\mathrm{T}}$ m/s was assumed. A temperature profile varying in the $x$ direction was prescribed as $T(x)=9000.0+2000\sin(2\pi x)$. A uniform $64\times 64$ grid was used for the discretization of the domain, and 3-rd degree polynomials were used in the DG method. All shock-capturing functions and positivity-preserving limiters were turned off, and the flux (38)–(41) was also used for the interface fluxes (instead of a Riemann solver), to ensure that no stabilization or dissipation schemes can impact the entropy conservation. The total entropy production rate (over all cells in the simulation domain), which can be defined as
$s_{t,\mathrm{tot}}(t)=\sum\frac{\partial s}{\partial\mathbf{u}}\cdot\frac{\partial\mathbf{u}}{\partial t}$ (58)
was computed via analysis callbacks available in Trixi.jl. Different values of the tolerance $r$ (as used in Alg. 2) were used, ranging from $0.5$ K to $10^{-6}$ K.
@@
Figure 5: Time-averaged absolute value of the entropy production rate as a function of the parameter $r$.
Figure 4 shows the total entropy production rate plotted over time for different values of the tolerance $r$. It can be clearly seen that the entropy is not conserved to machine precision, due to the errors in the computation of temperature and the integral part of the entropy; however, the overall magnitude of the entropy production rate is still very small. Moreover, lower values of $r$ lead to noticeably lower values of the entropy production rate. To better assess the impact of the value of $r$ on the entropy production rate, we plot the time-averaged absolute value of the entropy production rate as a function of $r$, as shown on Fig. 5. It can be seen that decreasing $r$ up to a value of $10^{-4}$ K leads to decreasing entropy production rates due to errors in the flux function; use of values of $r$ smaller than $10^{-4}$ K however does not lead to lower entropy production rates due to other sources of error becoming dominant.
@@ 4.3 Blast wave
Here $\phi=\tan^{-1}(y/x)$. The initial conditions of the problem correspond to a circular-shaped region with a radius of 0.5 m of higher density and pressure with a constant outward radial velocity. As in the previous case, the gas was assumed to be molecular nitrogen, with a vibrational spectrum described by the infinite harmonic oscillator model. A square domain of size $[-2,2]\times[-2,2]$ was used, discretized by a uniform $64\times 64$ grid. 3-rd degree polynomials were used in the DG method. Similar to the previous case, different values of the tolerance $r$ (as used in Alg. 2) were used, ranging from $0.5$ K to $10^{-6}$ K.
@@
Figure 9: Time-averaged absolute value of the entropy production rate as a function of the parameter $r$.
Figure 8 shows the total entropy production rate plotted over time for various values of the tolerance $r$ for a temperature discretization step $\Delta T$ of 1 K. Similar to the previous test case, it can be seen that whilst there is an error in the entropy production rate at higher values of $r$, it decreases rapidly when smaller values of $r$ are used. Figure 9 shows the time-averaged absolute value of the entropy production rate as a function of $r$. The error is almost of the order of machine precision when $r$ is taken equal to 10-6 K.
@@ 4.4 Flow around a cylinder
Finally, we consider a two-dimensional flow around a cylinder as an example of application of the developed flux function to a high-enthalpy flow with strong shocks. The free-stream parameters were chosen to be $v_{\infty}=5956$ m/s, $p_{\infty}=476$ Pa, $T_{\infty}=901$ K, based on the experimental HEG cylinder test-case [30]; the cylinder diameter is 0.045 m. The gas was assumed to be either molecular nitrogen with a molecular mass of $m=4.6517\times 10^{-26}$ kg or molecular oxygen with a molecular mass $m=5.3134\times 10^{-26}$ kg. The flow speed is approximately Mach 10.5, depending on the chosen model for the internal energy. Three models for the internal energy were considered:
@@
    Calorically perfect molecular nitrogen flow with $\gamma=1.4$.
@@
    Molecular oxygen flow with the infinite harmonic oscillator model for the vibrational spectrum with $\theta_{v}=2273.5$ K.
  3. 3.
    Molecular oxygen flow with the cut-off anharmonic oscillator model for the vibrational spectrum with $\theta_{v}=2273.5$ K and $\theta_{v,\mathrm{anh}}=17.366$ K, with the cut-off at the dissociation energy of oxygen $D=59364$ K. This corresponds to 36 vibrational levels accounted for.
For the energy and specific heat tables, a step size of $\Delta T=1$ K was used, and the tolerance $r$ in algorithm 2 was taken to be $10^{-5}\cdot\Delta T$.
@@
The simulations were performed on simple shock-fitted structured $60\times 60$ grids. A single cubic curve was fitted to the shock shape. As it does not fit the shock perfectly, some minor errors in the density and pressure are expected near certain regions of the shock, where the mesh and the shock are not well-aligned. Figure 10 shows an example of the simulation grid, specifically, the grid used for simulations with the infinite harmonic oscillator model. The origin of the cylinder is located at $(x,y)=(0,0)$. In the analysis of the simulations presented further below, flow quantities are considered along the stagnation line, given by $y=0$, and the outflow boundary, given by $x=0$. The following boundary conditions were applied in the simulation: supersonic inflow at the inflow boundary (highlighted in purple), supersonic outflow for the outflow boundary (highlighted in orange), and slip conditions at the stagnation line (highlighted in red) and cylinder surface (highlighted in blue).
The developed entropy-conserving flux was used for the volume fluxes, and the dissipative flux (57) was used for the surface fluxes in order to provide sufficient numerical diffusion for the strong shocks present in the flow. Due to the presence of strong shocks in the flow, the sub-cell shock capturing method of Hennemann et al. [24] was used.
The DLR TAU solver [36, 20] was used as a benchmark solver for this case with a similar grid and parameters governing the gas properties. The AUSM+ flux [35] was used, along with local time-stepping and a first-order implicit Backward Euler solver.
Firstly, we consider the simplest case, that of the calorically perfect molecular nitrogen flow with $\gamma=1.4$. This case has been chosen to verify the basic solver components in the Trixi.jl implementation, such as handling of curvilinear structured meshes, boundary conditions, and shock capturing.
Figure 11: Flow-field pressure computed the developed entropy-conservative approach implemented in Trixi.jl (top) and using the TAU code (bottom) for the case of a constant specific heat ($\gamma=1.4$).
Figure 11 shows the computed flow-field pressure for TAU and Trixi.jl simulations grid, with the Trixi.jl results computed using 2-nd order polynomials. Excellent agreement in the shock position and overall pressure profiles can be observed between the solvers. The Trixi.jl solution used a grid with several larger cells in the pre-shock region, therefore inflow region is somewhat larger on the top subplot. For a more quantitative comparison, we evaluate the flow properties along the stagnation line.
@@
Figure 13: Flow-field pressure computed the developed entropy-conservative approach implemented in Trixi.jl (top) and using the TAU code (bottom) for the case of the infinite harmonic oscillator.
Now we consider the cases with temperature-dependent specific heats. Figure 13 shows the computed flow-field pressure for the case of the vibrational spectrum modelled by an infinite harmonic oscillator. Similar to the previous case, excellent qualitative agreement can be seen between the simulations in the whole domain.
@@
For a more detailed analysis of this case and that of the cut-off anharmonic oscillator, we look at the flow-field properties along the stagnation line $y=0$ (assuming that the center of the cylinder is located at $(0,0)$). Figure 14 shows the pressure and temperature along the stagnation line (the surface of the cylinder is located at $x=-4.5$ cm) computed using TAU and Trixi.jl for the two different vibrational spectrum models, i.e. infinite harmonic oscillator and the cut-off anharmonic oscillator. Compared to the case of the calorically perfect gas with $\gamma=1.4$, both models for the vibrational spectrum have lower values of $\gamma$ (as seen on Fig. 1), thus leading to smaller shock stand-off distances [51]. The use of the cut-off harmonic oscillator model leads to smaller values of $\gamma$ compared to the infinite harmonic oscillator model, and therefore the stand-off distance is smaller for this case.
Finally, we analyse the flow quantities in the expansion region, namely, along the outflow boundary given by $x=0$. Slightly more differences were observed between Trixi.jl simulations using different grid and polynomial orders in this region, therefore, more results are presented: additional simulations were carried out for 2-nd, 3-rd, and 4-th order polynomials on $30\times 30$ and $60\times 60$ grids.
@@
Figure 15 shows the pressure and temperature in the expansion region along the line $x=0$ (the surface of the cylinder is located at $y=4.5$ cm) for the infinite harmonic oscillator model of the vibrational spectrum. Solid lines correspond to the $60\times 60$ grid, dotted lines correspond to the $30\times 30$ grid. We see that all simulation results agree very well, with only the 2-nd order DG simulation on the coarse $30\times 30$ grid (dotted red line) deviating slightly in terms of pressure and temperature from the other results. The Trixi.jl-based DG simulations predict a sharper shock, but use of the higher-order polynomials also leads to the presence of irregularities (seen on the temperature profile at $y=8$ cm), as well as slightly stronger oscillations near the shock (seen on the temperature profile near $y=9.7$ cm).